OpenStudy (anonymous):

Three charges Q1, Q2, and Q3 are equal to 6.4*10^-19 C. They are on a horz. straight line. The distance between each charge is 60nm. The point P is 80nm below Q2. What is the Magnitude of the Electric field at P? I have tried over and over! The correct answer is 1.9*10^10 N/C but I have no Idea how to do it! Thanks!

OpenStudy (anonymous):



OpenStudy (anonymous):

\[E=k\frac{ Q }{ r^2 }\]

OpenStudy (anonymous):

you will have to add up all the contributions at that point from the all the charges

OpenStudy (anonymous):

You will only have to focus on the y direction, because due to the symmetry of the charges the x component's cancel out.

OpenStudy (anonymous):

ok, this is where I am getting stuck.. \[E =\frac{8.99*10^{9} *6.4 *10^{-19}}{ 10^{4} } + \frac{8.99*10^{9} *6.4 *10^{-19}}{ 10^{4} } + \frac{8.99*10^{9} *6.4 *10^{-19} }{ 80^{2} }\] Which equals 2.04972 × 10^-12. The answer in the back says 1.9*10^10 .. Way off...

OpenStudy (anonymous):

units. the distance is given in nm, you want meters

OpenStudy (anonymous):

you also need to only be taking the y component from the two side charges. you will have to use a trig function (or well i would)

OpenStudy (anonymous):